Proton accelerated from rest thru potential difference of 1.0 MV. Final speed?

1.A proton (q=e, mp=1.67 x 10-27kg) is accelerated from rest through a potential difference of 1.0 MV. What is the final speed?
Answer #1

using conservation of energy

qV = .5mv^2

q = charge on proton = 1.6 x 10^-19 C
.5mv^2 = 1.6 x 10^-13
so

we get v = 1.3843  x 10^7 m/s
Answer #2

eV = 1/2mv^2

v = √[ 2eV/m]

plug the values, V = 1.38*10^7 m/s
Answer #3
mp*1/2*v2=q*V=>v=√(2Vq/m)
m=1.67*10^-27kg
q= 1.6*10^-19 C
V=10^6
plug the value we get v˜1.4*10^7m/s
Answer #4

Charge q= 1.6 * 10 ^ -19 C

Mass m = 1.67 * 10 ^ -27 kg

Potential difference V = 1 Mv = 10 ^ 6 V

Final speed v = √ [ 2Vq / m]

Plug the values we get v = 1.384 * 10^ 7 m / s

In two significant figures v = 1.4 * 10 ^ 7 m / s
Answer #5
we have
conservation of energy
mp*1/2*v2=q*106 V
mp=1.67 x 10^-27kg
v=1.4*107 m/s (ans)